A wire 50.0 m long 려nd 2.00 mm in diameter is connected to a source with potential difference of 9.11 V, and the current is found to be 5.72 A. Assume a temperature of 20°C and using the table below, identify the metal out of which the wire is mad Resistivities and Temperature Coefficients of Resistivity for Various Materials (at 20°c) Material Resistivity Temperature Coefficient (Q m)of Resistivity (°c)-1 -3 1.59× 10 3.8 x 10 1.7 x 10-8 2.44 x 10- 2.82 x 10 Copper 3.9 x 10 Gold 3.4 x 103 Aluminum 3.9 x 10 5.6 x 10-8 10.0× 10 11 x 10 4.5 x 10-3 5.0 x 10 3.92 x 10 Iron Lead 22 × 10 3.9× 10 -3 Nichrome150 x 108 0.4 x 10 Carbon -0.5 x 1043 3.5 x 10 0.46 -48 x 10 640 -75 x 103 1010-1014 Glass Hard rubber Sulfur 13 10 15 10 Quartz (fused 75x 1016 A nickel-chromium alloy commonly used in heating elements. O copper O gold O iron o silver
Answer #1

Solution) V= 9.11V

i= 5.72 A

L=50m

d=2mm

Radius r = d/2 =1mm=1×10^(-3)m

V=iR

R=V/i = 9.11/5.72 = 1.59 ohms

R = (resistivity)(L)/(A)

Resistivity = RA/L = (1.59)((pi)×(1×10^(-3))^2)/50

Resistivity = 1×10^(-7)=10×10^(-8) ohm meter

From the given table it should be iron

Therefore from given options , option c is correct

Iron is the answer